Using this equation (-6x +3y=2) determine the intercepts, write the equation in slope-intercept form, and write equations of a parallel and perpendicular line.
Thank you in advance.

Answer by checkley77(12844)

-6x+3y=2
3y=6x+2
y=6x/3+2/3
y=2x+2/3 y intercept=(0,2/3).
Set y=0 & solve for the x intercept.
0=2x+2/3
2x=-2/3
x=(-2/3)/2
x=-2/3*1/2
x=-2/6
x=-1/3 or (-1/3,0) the x intercept.
A parallel line will have the same slope:
y=2x+9 ans. for the parallel line.
y=-x/2-5 ans. for the perpendicular line.